OpenStudy (anonymous):

if y varies directly as x and y = 5 when x = 3 find y when x= 18

OpenStudy (anonymous):

y varies as x so y=kx find k by plugging in values of y as 5 and x as 3 then find y using value of k obtained from above and x as 18

OpenStudy (anonymous):

x/y=3/5=18/y Solve the proportion